We have a solvent called S316 with a density of 1.75 and we want to make 100 cc with a concentration of 50 ppm of an oil with a density of 0.8200 in S316 solvent. Please write its stoichiometric calculations
1
Expert's answer
2020-10-29T08:20:52-0400

50 ppm is equivalent to 50mg/L

And since we need 100cc

(1000/100)50=5mg=0.005g

Total mass =volume×density=100×0.82=820g

The mass of solvent= 820-0.005=819.995g

To find the volume of S316 solvent that will be needed,

V=mass/density

=819.995/1.75

=468.57ml